OpenStudy (anonymous):

Differentiate function f(x)= [(1/2)x]^5 the outside power, is messing me up

OpenStudy (anonymous):

\[(\frac{1}{2}x)^5\]

OpenStudy (anonymous):

if that's true \[\frac{1}{32}*5x^4\]

OpenStudy (anonymous):

so 5x^4/32?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks you soo much

OpenStudy (anonymous):

To see what is really going on under the hood you can apply the chain rule here, and make a substitution by setting u=(1/2)x\[\rightarrow\]du/dx=1/2 The equation can then be rewritten as: \[y=u^{5}\]\[\rightarrow\] dy/du=\[5u ^{4}\] From the chain rule dy/dx=dy/du * du/dx = (\[5u ^{4} * 1/2\] Since u = (1/2)x then dy/dx =\[ 5 (1/16)x ^{4} * 1/2\] = \[(5/32)x ^{4}\]